1. In an experiment, 9.80 g of olive oil at 80.0C are quickly transferred to water in a calorimeter at 18.0°C. The heat capacity of the water is 4.18 J/gC. The final equilibrium temperature is 22.8°C. What is the approximate specific heat capacity of the olive oil? 5. ) Given the standard enthalpies of formation for the following substances, determine the change in enthalpy for the combustion of 1.0 mol of propane, C3H8. Substance ΔH°f in kJ/mol C3H8(g) 103.8 CO2(g) 393.5 H2O(l) 285.8 6. Calculate the energy released as heat when 45.8 g of Fe(s) is combined with 27.9g of O2(g) if the ΔH°rxn is -1648 kJ/mol. The balanced equation for the reaction is 4 Fe(s) + 3O2(g) = 2 Fe2O3(s)
1
Expert's answer
2019-12-16T04:53:10-0500

1. Heat gained = Heat lost

"m_{water}s_{water}\Delta T = m_{olive}s_{olive}\Delta T"

"m*4.18*4.8=9.8*s_{oil}*(80-22.8)"

"\implies s_{oil}=m*2.047/57.2"

"=m*0.0358J\/g^\u00b0C"

where m is mass (in grams) of water in calorimeter.


5. Combustion of 1 mol propane takes place according to the following chemical reaction:

"C_3H_8(g)+ 5O_2(g) \to 3CO_2(g)+ 4H_2O(l)"

Thus change in enthalpy for this reaction is:

"\Delta H= H_{f(products)}-H_{f(reactants)}"

"=[3H_{f(CO_2\u200b)\u200b}+4H_{f(H_2\u200bO)}\u200b]\u2212""[H_{f(C_3\u200bH_8\u200b)\u200b}+5H_{f(O_2\u200b)}\u200b]"

"= 3*(-393.5)+ 4*(-285.8)-(-103.8)-5*0"

"\implies \Delta H= -2219.9 kJ"

Thus, heat released in this reaction is "2219.9 kJ" (Answer)


6. Given "\Delta H_{rxn} = -1648 kJ"

for 4 moles of "Fe(s)"

Thus heat released on combustion of 1 mol or 55.8g "Fe(s)" will be "1648\/4=412kJ"

Thus, heat released on combustion of 45.8g "Fe(s)" is "412*45.8\/55.8=35.44kJ" (Answer)